# Latent Factor Analysis in Short Panels

### v2

@@ 3 Feasible asymptotic distributional theory
In the proof of Proposition 2 below, we establish an asymptotic expansion for the LR test statistic under the null hypothesis of $k$ latent factors, by deriving an asymptotic expansion for $n\|\hat{S}\|^{2}$. For a $(T-k)\times(T-k)$ symmetric matrix $Z=(z_{i,j})$ and $p=\frac{1}{2}(T-k)(T-k+1)$, let us define the $p$-dimensional vector $vech(Z)=\left(\frac{1}{\sqrt{2}}z_{11},...,\frac{1}{\sqrt{2}}z_{T-k,T-k},\{z_{i,j}\}_{i<j}\right)^{\prime}$, where the pairs of indices $(i,j)$ with $i<j$ are ranked as $(1,2),(1,3),...,(1,T-k),(2,3),...,(T-k-1,T-k)$. [^13] This definition of the half-vectorization operator for symmetric matrices differs from the usual one by the ordering of the elements, and the rescaling of the diagonal elements. It is more convenient for our purposes (see proof of Lemma 10). For instance, it holds $\frac{1}{2}\|A\|^{2}=vech(A)^{\prime}vech(A)$, for a symmetric matrix $A$. Moreover, let us define the $p\times T$ matrix $\boldsymbol{X}=\left[vech(G^{\prime}E_{1,1}G)~{}:~{}\cdots~{}:~{}vech(G^{\prime}E_{T,T}G)\right]$, where $E_{t,t}$ denotes the $T\times T$ matrix with entry 1 in position $(t,t)$ and 00 elsewhere, and $G$ is a $T\times(T-k)$ matrix such that $F^{\prime}V_{\varepsilon}^{-1}G=0$ and $G^{\prime}V_{\varepsilon}^{-1}G=I_{T-k}$. Matrix $G$ is unique up to post-multiplication by an orthogonal matrix, and the columns of $\boldsymbol{X}$ span the linear space $\{vech(G^{\prime}DG):D\text{ diagonal}\}$. Then, the asymptotic expansion of $n\|\hat{S}\|^{2}$ under $H_{0}(k)$ is:
$\frac{n}{2}\|\hat{S}\|^{2}=vech(Z_{n}^{*})^{\prime}M_{\boldsymbol{X}}vech(Z_{n}^{*})+o_{p}(1),$ (5)
@@
where $Z^{*}_{n}:=G^{\prime}V_{\varepsilon}^{-1}Z_{n}V_{\varepsilon}^{-1}G$, $Z_{n}:=\sqrt{n}\left(\frac{1}{n}\varepsilon\varepsilon^{\prime}-\tilde{V}_{\varepsilon}\right)$, and matrix $M_{\boldsymbol{X}}:=I_{p}-\boldsymbol{X}(\boldsymbol{X}^{\prime}\boldsymbol{X})^{-1}\boldsymbol{X}^{\prime}$ is idempotent of rank $p-T=df$ [^14] We can extend results like (5) to test statistics that are generic functions of the eigenvalues of matrix $\hat{S}$ by using the Weyl’s inequalities (see e.g. Bernstein (2009)), and develop test statistics along the lines of FGS. The full-rank condition for matrix $\boldsymbol{X}$ corresponds to the local identification condition in Assumption A.5 ((see Lemma 6)), analogously as in linear regression. The expansion in (5) does not depend on the diagonal elements of $Z_{n}$, since $vech(G^{\prime}V_{\varepsilon}^{-1}diag(Z_{n})V_{\varepsilon}^{-1}G)$ is spanned by the columns of $\boldsymbol{X}$, and thus is annihilated by the projection matrix $M_{\boldsymbol{X}}$. Such an irrelevance of the diagonal elements of $Z_{n}$ is an implication of Proposition 1 (c).
@@
We now establish the distributional convergence $Z^{*}_{n}\Rightarrow Z^{*}$ as $n\rightarrow\infty$ and $T$ is fixed, where $Z^{*}$ is a Gaussian symmetric matrix variate. By the block structure in Assumption 2 (b), we can write $Z^{*}_{n}$ as a sum of independent zero-mean terms: $Z^{*}_{n}=\frac{1}{\sqrt{n}}\sum_{m=1}^{J_{n}}z^{*}_{m,n},$ where the variables in the triangular array $z^{*}_{m,n}=\sum_{i\in I_{m}}G^{\prime}V_{\varepsilon}^{-1}\left(\varepsilon_{i}\varepsilon_{i}^{\prime}-E[\varepsilon_{i}\varepsilon_{i}^{\prime}]\right)V_{\varepsilon}^{-1}G$ are independent across $m$ and such that $E[z^{*}_{m,n}]=0$. In Appendix B, we invoke the CLT for independent heterogeneous variables to $vech(Z^{*}_{n})=\frac{1}{\sqrt{n}}\sum_{m=1}^{J_{n}}vech(z^{*}_{m,n})$ and use Assumptions 2 (c) and (d) to check the Liapunov condition. We get $Z^{*}_{n}\Rightarrow Z^{*}$, where $vech(Z^{*})\sim N(0,\Omega_{Z^{*}})$ and $\Omega_{Z^{*}}=\underset{n\rightarrow\infty}{\lim}\frac{1}{n}\sum_{m=1}^{J_{n}}V[vech(z^{*}_{m,n})]$. Then, the asymptotic expansion in (5) yields the asymptotic distribution for the LR test statistic defined in (4) under the null hypothesis, and is given in Proposition 2 (a) below.
@@
In Proposition 2 (b) hereafter, we use $\hat{G}=\hat{V}_{\varepsilon}^{1/2}\hat{Q}$, where $\hat{Q}$ is the $T\times(T-k)$ matrix of standardized eigenvectors of $\hat{S}$ corresponding to its $T-k$ non-zero eigenvalues ranked in decreasing order (see Proposition 1 (a)),151515 $\hat{Q}$ is the $T\times(T-k)$ matrix of standardized eigenvectors of $\hat{V}^{-1/2}_{\varepsilon}\hat{V}_{y}\hat{V}^{-1/2}_{\varepsilon}$ corresponding to its $T-k$ smallest eigenvalues ranked in decreasing order. and $\hat{\boldsymbol{X}}$ the matrix obtained by replacing $G$ with $\hat{G}$ in $\boldsymbol{X}$.
@@
[…]
Proposition 2 (a) shows that we depart from classical FA theory since we have convergence to a weighted sum of chi-square variates instead of a single chi-square variate (see the discussion in the next section). In Proposition 2 (b), matrix $\hat{G}$ is consistent for $G$ up to a rotation. The eigenvalues $\hat{\mu}_{j}$ are unaffected by such rotation because eigenvalues are invariant under pre and post multiplication by an orthogonal matrix and its transpose. [^16] In Proposition 9 in Appendix E.6, we study how $\boldsymbol{X}$ and $M_{\boldsymbol{X}}\Omega_{Z^{*}}M_{\boldsymbol{X}}$ are transformed under different choices for the rotation of $G$. The eigenvalues $\mu_{j}$ are invariant to such rotation as expected. With fixed $T$, the GLS residuals $\hat{\varepsilon}_{i}$ are asymptotically close to $M_{F,V_{\varepsilon}}\varepsilon_{i}$ and not to the true errors $\varepsilon_{i}$. However, it does not impede the consistency of the eigenvalues $\hat{\mu}_{j}$, since $G^{\prime}V_{\varepsilon}^{-1}M_{F,V_{\varepsilon}}=G^{\prime}V_{\varepsilon}^{-1}$. When we apply the CLT, centering is done implicitly since $M_{\boldsymbol{X}}vech(G^{\prime}V_{\varepsilon}^{-1}\varepsilon_{i}\varepsilon_{i}^{\prime}V_{\varepsilon}^{-1}G)=M_{\boldsymbol{X}}vech(G^{\prime}V_{\varepsilon}^{-1}(\varepsilon_{i}\varepsilon_{i}^{\prime}-E[\varepsilon_{i}\varepsilon_{i}^{\prime}])V_{\varepsilon}^{-1}G)$. Indeed $vech(G^{\prime}V_{\varepsilon}^{-1}E[\varepsilon_{i}\varepsilon_{i}^{\prime}]V_{\varepsilon}^{-1}G)$ is spanned by the columns of $\boldsymbol{X}$ since $E[\varepsilon_{i}\varepsilon_{i}^{\prime}]=\sigma_{ii}V_{\varepsilon}$ is diagonal. We can consistently estimate the critical values of the statistics by simulating a large number of draws from $\sum_{j=1}^{df}\hat{\mu}_{j}\chi^{2}_{j}(1)$. Hence, even if FA residuals are not consistent estimates of the true errors in short panels, we are still able to supply a feasible asymptotic distributional theory for our empirical applications under a block-dependence structure. Proposition 2 (c) gives test consistency against global alternatives.
@@ 4 Impact of assuming Gaussian errors and sphericity
In this section, we particularize the general distributional results of Proposition 2 to three important cases, namely Gaussian errors, settings where the asymptotic distribution under Gaussian errors still holds for the LR test statistics (up to scaling), and spherical errors. We further discuss the impact of wrongly assuming Gaussian errors or sphericity when conducting inference on the number of latent factors.
@@ 4.1 Gaussian errors
Let us consider the case where the errors $\varepsilon_{i}\overset{ind}{\sim}N(0,\sigma_{ii}V_{\varepsilon})$ are independent Gaussian vectors. From classical FA theory, we expect that the statistic $LR(k)$ admits asymptotically a chi-square distribution with $df$ degrees of freedom in the cross-sectionally homoskedastic case, i.e., $\sigma_{ii}=1$ for all assets $i$. We cannot expect that this distributional result applies to the Gaussian framework in full generality, since - even in such a case - our setting corresponds to a pseudo model (because the $\sigma_{ii}$ may be heterogeneous across $i$, and the $\beta_{i}$ are treated as fixed effects, namely incidental parameters, instead of Gaussian random effects). Under the normality assumption for the error terms, we have ${\varepsilon}^{*}_{i}:=G^{\prime}V_{\varepsilon}^{-1}\varepsilon_{i}\overset{ind}{\sim}N(0,\sigma_{ii}I_{T-k})$. Thus, by the Liapunov CLT, the distributional limit of $\frac{1}{\sqrt{q}}Z_{n}^{*}=\sqrt{n/q}\left(\frac{1}{n}{\varepsilon^{*}}(\varepsilon^{*})^{\prime}-\frac{1}{n}E\left[{\varepsilon^{*}}(\varepsilon^{*})^{\prime}\right]\right)$ is in the Gaussian Orthogonal Ensemble (GOE) for dimension $T-k$ (see e.g. Tao (2012)), i.e., $\frac{1}{\sqrt{q}}vech(Z^{*})\sim N(0,I_{p})$, where $q:=\underset{n\rightarrow\infty}{\lim}\frac{1}{n}\sum_{i=1}^{n}\sigma_{ii}^{2}$. Then, from (5) we get $LR(k)\Rightarrow q\chi^{2}(df)$, i.e., we get convergence to a scaled chi-square variate $q\chi^{2}(df)$. In the cross-sectionally homoskedastic case, we have $q=1$ yielding the classical $\chi^{2}(df)$ result. On the contrary, cross-sectional heterogeneity in the unconditional idiosyncratic variances yields $q>1$ and a deviation from classical FA theory even in the Gaussian case. In the Gaussian case, unobserved heterogeneity across asset idiosyncratic variances leads to an oversized LR test if we use critical values from the chi-square table without proper scaling. In our Monte Carlo experiments with non-Gaussian errors, we get size distortion by around 80 percentage points when $T=24$ for a nominal size of 5%, and thus an average estimated number of factors over 3 instead of 2.
@@ 4.2 Validity of the scaled asymptotic chi-square test
Let Assumptions 1-2, A.1-A.6 hold with (a) $\underset{n\rightarrow\infty}{\lim}\frac{1}{n}\sum_{i=1}^{n}E[\varepsilon_{i,t}\varepsilon_{i,s}\varepsilon_{i,r}\varepsilon_{i,p}]=qV_{\varepsilon,tt}V_{\varepsilon,ss}$, when $t=r>s=p$, for a constant $q>0$, and $=0$ in all other cases with $t>s$ and $r>p$, and (b) let $\kappa=\underset{n\rightarrow\infty}{\lim}\frac{1}{n}\sum_{m=1}^{J_{n}}\sum_{i\neq j\in I_{m}}\sigma_{ij}^{2}$ as in Assumption A.3 (b). Then, $LR(k)\Rightarrow\bar{q}\chi^{2}(df)$ under $H_{0}(k)$ for $\bar{q}:=q+\kappa$.
Conditions (a) and (b) in Proposition 3 generalize the correctness of the scaled chi-square test beyond Gaussianity and error independence across time and assets. Under Assumption 2, Condition (a) is satisfied if the standardized error terms $w_{i,t}$ are conditionally homoskedastic martingale difference sequences. However, Condition (a) excludes empirically relevant cases such as ARCH processes for $w_{i,t}$, because, in that case, $\frac{1}{V_{\varepsilon,tt}V_{\varepsilon,ss}}E[\varepsilon_{i,t}^{2}\varepsilon_{i,s}^{2}]$ depends on lag $t-s$. Hence, serial correlation in squared idiosyncratic errors is responsible for the deviation of the LR test from the scaled chi-square asymptotic distribution. This setting is covered by the general results in Proposition 2. Anderson and Amemiya (1988) establish the asymptotic distribution of FA estimates assuming that the error terms are i.i.d. across sample units and deploy an assumption that is analogue to Condition (a) above in their Corollary 2. The i.i.d. assumption in our case implies $\sigma_{ii}=1$ for all $i$, which results in a cross-sectionally homoskedastic setting. [^17] If the $\sigma_{ii}$ are treated as i.i.d. random effects independent of the errors, and we exclude cross-sectional correlation of errors to simplify, we recover the i.i.d. condition of the data. However, the random $\sigma_{ii}$ yields a stochastic common factor across time that breaks the condition in Corollary 2 of Anderson and Amemiya (1988). That setting is irrealistic in our application, as it would imply that the idiosyncratic variance is the same for all assets. Our results show that establishing the asymptotic distribution of the test statistics, especially the AUMPI property of LR test (see Section 5), in a general setting with non-Gaussian errors, heterogeneous idiosyncratic variances and ARCH effects, is challenging, but still possible.

### v3

@@ 4 Feasible asymptotic distributional theory
In the proof of Proposition 1 below, we establish an asymptotic expansion for the LR test statistic under the null hypothesis of $k$ latent factors. For integer $p=\frac{1}{2}(T-k)(T-k+1)$, let us define the $p\times T$ matrix $\boldsymbol{X}=\left[vech(G^{\prime}E_{1,1}G)~:~\cdots~:~vech(G^{\prime}E_{T,T}G)\right]$, where $E_{t,t}$ denotes the $T\times T$ matrix with entry 1 in position $(t,t)$ and 0 elsewhere, and $G$ is a $T\times(T-k)$ matrix such that $F^{\prime}V_{\varepsilon}^{-1}G=0$ and $G^{\prime}V_{\varepsilon}^{-1}G=I_{T-k}$. Matrix $G$ yields the orthogonal complement to the factor space in the scalar product induced by the inverse error variance, and is unique up to post-multiplication by an orthogonal matrix. The columns of $\boldsymbol{X}$ span the linear space $\{vech(G^{\prime}DG):D\text{ diagonal}\}$. Then, the asymptotic expansion of $LR(k)$ under $H_{0}(k)$ is:
$LR(k)=vech(Z_{n}^{*})^{\prime}M_{\boldsymbol{X}}vech(Z_{n}^{*})+o_{p}(1),$ (3)
@@
where $Z^{*}_{n}:=G^{\prime}V_{\varepsilon}^{-1}Z_{n}V_{\varepsilon}^{-1}G$ for $Z_{n}:=\sqrt{n}\left(\frac{1}{n}\varepsilon\varepsilon^{\prime}-\tilde{V}_{\varepsilon}\right)$, and matrix $M_{\boldsymbol{X}}:=I_{p}-\boldsymbol{X}(\boldsymbol{X}^{\prime}\boldsymbol{X})^{-1}\boldsymbol{X}^{\prime}$ is idempotent of rank pT=dfp-T=df. The full-rank condition for matrix $\boldsymbol{X}$ corresponds to the local identification condition for the PML estimator in Assumption 7 (see Lemma 7), analogously as in linear regression. The symmetric random matrix $Z_{n}$ involves average squares and cross-moments of error terms. The expansion in (3) does not depend on the diagonal elements of $Z_{n}$, since $vech(G^{\prime}V_{\varepsilon}^{-1}diag(Z_{n})V_{\varepsilon}^{-1}G)$ is spanned by the columns of $\boldsymbol{X}$, and thus is annihilated by the projection matrix $M_{\boldsymbol{X}}$. Intuitively, the diagonal elements of $Z_{n}$ are irrelevant because only the cross-sectional correlations of residuals for two different dates are useful to check for omitted factors.
@@
We now outline the distributional convergence $Z^{*}_{n}\Rightarrow Z^{*}$ as $n\rightarrow\infty$ and $T$ is fixed, where $Z^{*}$ is a Gaussian symmetric matrix variate. By the block structure in Assumption 2 (b), we can write $Z^{*}_{n}$ as a sum of independent zero-mean terms: $Z^{*}_{n}=\frac{1}{\sqrt{n}}\sum_{m=1}^{J_{n}}z^{*}_{m,n},$ where the variables in the triangular array $z^{*}_{m,n}=\sum_{i\in I_{m}}G^{\prime}V_{\varepsilon}^{-1}\left(\varepsilon_{i}\varepsilon_{i}^{\prime}-E[\varepsilon_{i}\varepsilon_{i}^{\prime}]\right)V_{\varepsilon}^{-1}G$ are independent across $m$ and such that $E[z^{*}_{m,n}]=0$. In Appendix A, we invoke the CLT for independent heterogeneous variables to $vech(Z^{*}_{n})=\frac{1}{\sqrt{n}}\sum_{m=1}^{J_{n}}vech(z^{*}_{m,n})$ and use Assumptions 2 (c) and (d) to check the Liapunov condition. We get $Z^{*}_{n}\Rightarrow Z^{*}$, where $vech(Z^{*})\sim N(0,\Omega_{Z^{*}})$ and $\Omega_{Z^{*}}=\underset{n\rightarrow\infty}{\lim}\frac{1}{n}\sum_{m=1}^{J_{n}}V[vech(z^{*}_{m,n})]$. Then, the asymptotic expansion in (3) yields the asymptotic distribution for the LR test statistic defined in (2) under the null hypothesis, and is given in Proposition 1 (a) below.
@@
In Proposition 1 (b) hereafter, we state a feasible result for the asymptotic distribution of the LR statistic. We use the estimate $\hat{G}$, i.e. the $T\times(T-k)$ matrix of standardized eigenvectors of $\hat{V}_{y}\hat{V}^{-1}_{\varepsilon}$ corresponding to its TkT-k smallest eigenvalues ranked in decreasing order such that $\hat{G}^{\prime}\hat{V}^{-1}_{\varepsilon}\hat{G}=I_{T-k}$, and set $\hat{\boldsymbol{X}}$ the matrix obtained by replacing $G$ with $\hat{G}$ in $\boldsymbol{X}$. The GLS projection matrix orthogonal to $F$ for variance $V$ is defined by $M_{F,V}:=I_{T}-F(F^{\prime}V^{-1}F)^{-1}F^{\prime}V^{-1}$.
@@
Let Assumptions 1-9 hold. As $n\rightarrow\infty$ and $T$ is fixed, under the null hypothesis $H_{0}(k)$ of $k$ latent factors, (a) $LR(k)\Rightarrow\sum_{j=1}^{df}\mu_{j}\chi^{2}_{j}(1)$, where the $\chi^{2}_{j}(1)$ are independent chi-square variables with one degree of freedom, and the $\mu_{j}$ are the $df$ non-zero eigenvalues of matrix $M_{\boldsymbol{X}}\Omega_{Z^{*}}M_{\boldsymbol{X}}$. (b) $\hat{\mu}_{j}\overset{p}{\rightarrow}\mu_{j}$, where $\hat{\mu}_{j}$, j=1,,dfj=1,...,df, are the non-zero eigenvalues of matrix $M_{\hat{\boldsymbol{X}}}\hat{\Omega}_{Z^{*}}M_{\hat{\boldsymbol{X}}}$, and $\hat{\Omega}_{Z^{*}}=\frac{1}{n}\sum_{m=1}^{J_{n}}vech(\hat{z}^{*}_{m,n})vech(\hat{z}^{*}_{m,n})^{\prime}$ with $\hat{z}^{*}_{m,n}=\sum_{i\in I_{m}}\hat{G}^{\prime}\hat{V}_{\varepsilon}^{-1}\hat{\varepsilon}_{i}\hat{\varepsilon}_{i}^{\prime}\hat{V}_{\varepsilon}^{-1}\hat{G}$ and $\hat{\varepsilon}_{i}=M_{\hat{F},\hat{V}_{\varepsilon}}(y_{i}-\bar{y})$. Under the alternative hypothesis $H_{1}(k)$ of more than $k$ latent factors, (c) $LR(k)\geq Cn$, w.p.a. $1$ for a constant $C>0$, and $\hat{\mu}_{j}=O_{p}(n\sum_{m=1}^{J_{n}}B_{m,n}^{2})=o_{p}(n)$.
Proposition 1 (a) shows that we depart from classical FA theory since we have convergence to a weighted sum of chi-square variates (see Section 4 of Robin and Smith (2000) for discussion of such random variable) instead of a single chi-square variate (see the discussion in Section 5). Specification testing in the GMM framework ($J$-test of Hansen (1982)) without an optimal weighting matrix also gives this type of asymptotic distribution. It differs from the distribution of a mixture of chi-squared distributions found in testing when a parameter is on the boundary of the maintained hypothesis (see e.g. Andrews (2001)). In Proposition 1 (b), matrix $\hat{G}$ is consistent for $G$ up to a right-rotation. The eigenvalues $\hat{\mu}_{j}$ are unaffected by such rotation because eigenvalues are invariant under pre and post multiplication by an orthogonal matrix and its transpose. In Proposition 8 in Appendix B.6, we study how $\boldsymbol{X}$ and $M_{\boldsymbol{X}}\Omega_{Z^{*}}M_{\boldsymbol{X}}$ are transformed under different choices for the rotation of $G$. The eigenvalues $\mu_{j}$ are invariant to such rotation as expected. With fixed $T$, the GLS residuals $\hat{\varepsilon}_{i}$ are asymptotically close to $M_{F,V_{\varepsilon}}\varepsilon_{i}$ and not to the true errors $\varepsilon_{i}$. However, it does not impede the consistency of the eigenvalues $\hat{\mu}_{j}$ underlying our feasible inference, since $G^{\prime}V_{\varepsilon}^{-1}M_{F,V_{\varepsilon}}=G^{\prime}V_{\varepsilon}^{-1}$. When we apply the CLT, centering is done implicitly since $M_{\boldsymbol{X}}vech(G^{\prime}V_{\varepsilon}^{-1}\varepsilon_{i}\varepsilon_{i}^{\prime}V_{\varepsilon}^{-1}G)=M_{\boldsymbol{X}}vech(G^{\prime}V_{\varepsilon}^{-1}(\varepsilon_{i}\varepsilon_{i}^{\prime}-E[\varepsilon_{i}\varepsilon_{i}^{\prime}])V_{\varepsilon}^{-1}G)$. Indeed $vech(G^{\prime}V_{\varepsilon}^{-1}E[\varepsilon_{i}\varepsilon_{i}^{\prime}]V_{\varepsilon}^{-1}G)$ is spanned by the columns of $\boldsymbol{X}$ since $E[\varepsilon_{i}\varepsilon_{i}^{\prime}]=\sigma_{ii}V_{\varepsilon}$ is diagonal. We can consistently estimate the critical values of the statistics by simulating a large number of draws of $\sum_{j=1}^{df}\hat{\mu}_{j}\chi^{2}_{j}(1)$. Hence, even if FA residuals are not consistent estimates of the true errors in short panels, we are still able to supply a feasible asymptotic distributional theory for our empirical applications under a block-dependence structure. Proposition 1 (c) gives test consistency against global alternatives (see Section 6 for local alternatives and AUMPI properties).
@@ 5 Impact of assuming Gaussian errors and sphericity
In this section, we particularize the general distributional results of Proposition 1 to three important cases, namely (i) Gaussian errors, (ii) settings where the asymptotic distribution under Gaussian errors still holds for the LR test statistics (up to scaling), and (iii) spherical errors. We further discuss the impact of wrongly assuming Gaussian errors or sphericity (PCA) when conducting inference on the number of latent factors. This section should help guard against misuse of standard software routines in applied work.
@@ 5.1 Gaussian errors
Let us consider the case where the errors $\varepsilon_{i}\overset{ind}{\sim}N(0,\sigma_{ii}V_{\varepsilon})$ are independent Gaussian vectors. From classical FA theory, we expect that the statistic $LR(k)$ admits asymptotically a chi-square distribution with $df$ degrees of freedom in the cross-sectionally homoskedastic case, i.e., $\sigma_{ii}=1$ for all assets $i$. We cannot expect that this distributional result applies to the Gaussian framework in full generality, since - even in such a case - our setting corresponds to a pseudo model (because the $\sigma_{ii}$ may be heterogeneous across $i$, and the $\beta_{i}$ are treated as fixed effects, namely non-random incidental parameters, instead of Gaussian random effects). Under the normality assumption for the error terms, we have ${\varepsilon}^{*}_{i}:=G^{\prime}V_{\varepsilon}^{-1}\varepsilon_{i}\overset{ind}{\sim}N(0,\sigma_{ii}I_{T-k})$. Thus, by the Liapunov CLT, the distributional limit of $\frac{1}{\sqrt{q}}Z_{n}^{*}=\sqrt{n/q}\left(\frac{1}{n}{\varepsilon^{*}}(\varepsilon^{*})^{\prime}-\frac{1}{n}E\left[{\varepsilon^{*}}(\varepsilon^{*})^{\prime}\right]\right)$ is in the Gaussian Orthogonal Ensemble (GOE) for dimension TkT-k (see e.g. Tao (2012)), i.e., $\frac{1}{\sqrt{q}}vech(Z^{*})\sim N(0,I_{p})$, where $q:=\underset{n\rightarrow\infty}{\lim}\frac{1}{n}\sum_{i=1}^{n}\sigma_{ii}^{2}$. Then, from (3) we get $LR(k)\Rightarrow q\chi^{2}(df)$, i.e., we get convergence to a scaled chi-square variate $q\chi^{2}(df)$. In the cross-sectionally homoskedastic case, we have $q=1$ yielding the classical $\chi^{2}(df)$ result. On the contrary, cross-sectional heterogeneity in the unconditional idiosyncratic variances yields $q>1$ and a deviation from classical FA theory even in the Gaussian case. In the Gaussian case, unobserved heterogeneity across asset idiosyncratic variances leads to an oversized LR test if we use critical values from the chi-square table without proper scaling. In our Monte Carlo experiments with non-Gaussian errors, we get size distortion by around 80 percentage points when $T=24$ for a nominal size of 5%, and thus an average estimated number of factors over 3 instead of 2.
@@ 5.2 Validity of the scaled asymptotic chi-square test
Let Assumptions 1-8 hold with (a) $\underset{n\rightarrow\infty}{\lim}\frac{1}{n}\sum_{i=1}^{n}E[\varepsilon_{i,t}\varepsilon_{i,s}\varepsilon_{i,r}\varepsilon_{i,p}]=qV_{\varepsilon,tt}V_{\varepsilon,ss}$, when $t=r>s=p$, for a constant $q>0$, and =0=0 in all other cases with $t>s$ and $r>p$, and (b) let $\kappa=\underset{n\rightarrow\infty}{\lim}\frac{1}{n}\sum_{m=1}^{J_{n}}\sum_{i\neq j\in I_{m}}\sigma_{ij}^{2}$ as in Assumption 5 (b). Then, $LR(k)\Rightarrow\bar{q}\chi^{2}(df)$ under $H_{0}(k)$ for $\bar{q}:=q+\kappa$.
Conditions (a) and (b) in Proposition 2 generalize the correctness of the scaled chi-square test beyond Gaussianity and error independence across time and assets. Under Assumption 2, Condition (a) is satisfied if the standardized error terms $w_{i,t}$ are conditionally homoskedastic martingale difference sequences. However, Condition (a) excludes empirically relevant cases such as ARCH processes for $w_{i,t}$, because, in that case, $\frac{1}{V_{\varepsilon,tt}V_{\varepsilon,ss}}E[\varepsilon_{i,t}^{2}\varepsilon_{i,s}^{2}]$ depends on lag tst-s. Hence, serial correlation in squared idiosyncratic errors is responsible for the deviation of the LR test from the scaled chi-square asymptotic distribution. This setting is covered by the general results in Proposition 1. Anderson and Amemiya (1988) establish the asymptotic distribution of FA estimates assuming that the error terms are i.i.d. across sample units and deploy an assumption that is analogue to Condition (a) above in their Corollary 2. The i.i.d. assumption in our case implies $\sigma_{ii}=1$ for all $i$, which results in a cross-sectionally homoskedastic setting. That setting is irrealistic in our application, as it would imply that the idiosyncratic variance is the same for all assets. Our results show that establishing the asymptotic distribution of the test statistics, especially the AUMPI property of LR test (see Section 6), in a general setting with non-Gaussian errors, heterogeneous idiosyncratic variances and ARCH effects (individual and common across assets), is challenging, but still possible. If the $\sigma_{ii}$ are treated as i.i.d. random effects independent of the errors, and we exclude cross-sectional correlation of errors to simplify, we recover the i.i.d. condition of the data. However, the random $\sigma_{ii}$ yields a stochastic common factor across time that breaks the condition in Corollary 2 of Anderson and Amemiya (1988).
When the assumptions of Proposition 2 are satisfied, scaling the LR test statistic by a consistent estimate of $1/\bar{q}$ is sufficient to recover an asymptotic $\chi^{2}(df)$ distribution, as in the Gaussian case. The idea of scaling a test statistic to make it more robust to deviations from Gaussianity is well known in the psychometrics literature, for example, in the Satorra-Bentler scaled difference chi-square test (Satorra and Bentler (2001, 2010)). In our context, the Satorra-Bentler scaling factor corresponds to a consistent estimate of $df/\sum_{j=1}^{df}\mu_{j}$, and ensures that the weighted sum of chi-square variates described in Proposition 1 has the same mean as a $\chi^{2}(df)$ variate. Under the conditions of Proposition 2, we have $\frac{1}{df}\sum_{j=1}^{df}\mu_{j}=\bar{q}$, so that $1/\bar{q}$ coincides with the Satorra-Bentler scaling factor. Unreported Monte Carlo experiments under the design of Section 7 show that this scaling approach performs well in finite samples.